OpenStudy (anonymous):

A fuel oil distributor has 120000 gallons of fuel with 0.9% sulfur content which exceeds pollution standards of 0.8%, sulfur content. How many gallons of fuel oil with a 0.3% sulfur content must be added to the 120000 gallons to obtain fuel oil that weight pollution content standards.

OpenStudy (jack1):

sorry i had those equations mixed up before: \[ \frac{(1080 + 0.003x)}{(120000 + x)} = 0.008 \] now solve for x \[ (1080 + 0.003x) = 0.008 \times (120000 + x) \] \[ 1080 + 0.003x = (0.008 \times 120000) + (0.008 \times x) \] \[ 1080 + 0.003x = 960 + 0.008 x \] \[ 1080 -960 = 0.008 x - 0.003x \] \[ 120 = 0.005 x \] x = ...?

OpenStudy (jack1):

note: 120000 gallons of fuel with 0.9% sulfur content = 1080 gal sulfur